OpenStudy (anonymous):

Raising any base to the power of zero will yield _____. A. 1 B. 0 C. –1 D. a nonreal number

OpenStudy (tkhunny):

If we remember that "any base" means any VALID base, then it is usually defined at Unity. 10^2 = 100 10^1 = 10 10^0 = 1 2^3 = 8 2^2 = 4 2^1 = 2 2^0 = 1

OpenStudy (anonymous):

thanks